OpenStudy (anonymous):

"cos(ω' T) is by definition equal to one"? Explanation wanted! I can't wrap my head around it. I know each T is supposed to be one period. So it will always be 1, 2, 3 and so on periods; which then the COS will be at the top (1); however the ω is throwing me off. :S

OpenStudy (michele_laino):

hint: by definition of period T, we have: \[\Large \omega T = \omega \cdot \frac{{2\pi }}{\omega } = 2\pi \]

OpenStudy (michele_laino):

therefore: \[\Large \cos \left( {\omega T} \right) = \cos \left( {2\pi } \right) = ...?\]

OpenStudy (anonymous):

Ah, thank you! Understand! Of course! xP

OpenStudy (michele_laino):

:)